# Test-time Correlation Alignment

### v1

@@ A C P R
Method $t\xrightarrow{\hskip 455.24408pt}$ Avg
@@
Table 11: Accuracy comparisons of different TTA methods on CIFAR-10-C dataset at damage level of 5, based on ResNet-18 backbone with 15 types of damage applied sequentially to a continuously adapted model. The best results are highlighted in boldface, and the second ones are underlined.
Method $t\xrightarrow{\hskip 455.24408pt}$ Avg
@@
Table 12: Accuracy comparisons of different TTA methods on CIFAR-10-C dataset at damage level of 5, based on ResNet-50 backbone with 15 types of damage applied sequentially to a continuously adapted model. The best results are highlighted in boldface, and the second ones are underlined.
Method $t\xrightarrow{\hskip 455.24408pt}$ Avg
@@
TENT (Wang et al., 2020) 36.60 43.79 39.41 83.28 70.44 83.72 85.94 87.19 85.21 79.33 92.76 29.43 84.73 84.90 84.62 71.42
@@
SAR (Niu et al., 2023) 37.23 44.30 39.93 83.16 70.32 83.54 85.80 87.15 85.06 79.19 92.75 29.75 84.73 84.68 84.58 71.48
TIPI (Nguyen et al., 2023) 36.46 43.79 39.30 83.30 70.45 83.68 85.89 87.18 85.20 79.29 92.74 29.27 84.69 84.91 84.59 71.38
TEA (Yuan et al., 2024) 36.59 43.83 39.30 83.36 70.43 83.71 85.96 87.17 85.18 79.48 92.74 29.83 84.72 84.87 84.59 71.45
@@
Table 13: Accuracy comparisons of different TTA methods on CIFAR-10-C dataset at damage level of 5, based on ViT-B/16 backbone with 15 types of damage applied sequentially to a continuously adapted model. The best results are highlighted in boldface, and the second ones are underlined.
Method $t\xrightarrow{\hskip 455.24408pt}$ Avg
@@
Table 14: Accuracy comparisons of different TTA methods on CIFAR-100-C dataset at damage level of 5, based on ResNet-18 backbone with 15 types of damage applied sequentially to a continuously adapted model. The best results are highlighted in boldface, and the second ones are underlined.
Method $t\xrightarrow{\hskip 455.24408pt}$ Avg
@@
Table 15: Accuracy comparisons of different TTA methods on CIFAR-100-C dataset at damage level of 5, based on ResNet-50 backbone with 15 types of damage applied sequentially to a continuously adapted model. The best results are highlighted in boldface, and the second ones are underlined.
Method $t\xrightarrow{\hskip 455.24408pt}$ Avg
@@
Table 16: Accuracy comparisons of different TTA methods on CIFAR-100-C dataset at damage level of 5, based on ViT-B/16 backbone with 15 types of damage applied sequentially to a continuously adapted model. The best results are highlighted in boldface, and the second ones are underlined.

### v2

@@ A C P R
Backbone Method DomainNet Avg Hyper-parameters
C I P Q R S
ResNet-18 Source (He et al., 2016) 57.30 16.86 45.03 12.69 56.89 46.00 39.13 nan
BN (Schneider et al., 2020) 57.26 11.55 43.32 11.77 56.58 47.09 37.93 nan
T3A (Iwasawa & Matsuo, 2021) 58.44 18.57 46.80 14.54 57.66 47.72 40.62 fk=100
AdaNPC (Zhang et al., 2023) 57.61 15.83 44.89 18.44 59.72 46.53 40.50 fk=100 k=10
TENT (Wang et al., 2020) 58.41 13.09 45.17 13.02 57.89 48.94 39.42 lr=1e-4
PLC (Lee, 2013) 57.45 12.60 44.77 10.01 55.74 47.20 37.96 lr=1e-5
EATA (Niu et al., 2022) 59.18 16.22 46.65 18.04 59.59 50.21 41.65 lr=1e-3
SAR (Niu et al., 2023) 59.13 13.10 45.75 4.88 58.25 49.83 38.49 lr=5e-3
TIPI (Nguyen et al., 2023) 58.42 11.68 42.53 5.37 50.76 47.58 36.05 lr=5e-4
TEA (Yuan et al., 2024) 58.01 12.83 45.10 14.33 57.55 48.80 39.43 lr=5e-5
TSD (Wang et al., 2023) 57.73 12.19 44.58 12.78 55.94 48.31 38.59 lr=1e-5 fk=100
[…]
Table 15: Accuracy comparison of different TTA methods on DomainNet dataset based on ResNet-18 backbone. The best results are highlighted in boldface, and the second ones are underlined.
Backbone Method DomainNet Avg Hyper-parameters
C I P Q R S
ResNet-50 Source (He et al., 2016) 63.68 20.93 50.35 12.95 62.16 51.42 43.58 nan
BN (Schneider et al., 2020) 63.30 14.84 48.54 10.83 62.02 52.12 41.94 nan
T3A (Iwasawa & Matsuo, 2021) 63.76 21.06 49.82 18.46 64.05 52.39 44.92 fk=100
AdaNPC (Zhang et al., 2023) 64.38 20.12 51.07 17.34 65.59 52.51 45.17 fk=200 k=10
TENT (Wang et al., 2020) 64.95 17.46 51.58 11.28 64.04 54.51 43.97 lr=1e-4
PLC (Lee, 2013) 63.56 14.89 49.09 8.83 60.83 52.61 41.63 lr=1e-5
EATA (Niu et al., 2022) 65.89 19.88 52.67 20.36 66.58 55.99 46.89 lr=5e-4
SAR (Niu et al., 2023) 65.44 14.63 50.68 3.94 63.94 54.49 42.19 lr=5e-3
TIPI (Nguyen et al., 2023) 64.97 17.47 51.63 11.67 64.03 54.69 44.08 lr=1e-4
TEA (Yuan et al., 2024) 64.87 16.95 51.40 11.48 61.90 54.31 43.48 lr=1e-4
TSD (Wang et al., 2023) 64.31 16.53 50.75 8.52 58.97 53.63 42.12 lr=5e-5 fk=5
[…]
Table 16: Accuracy comparison of different TTA methods on DomainNet dataset based on ResNet-50 backbone. The best results are highlighted in boldface, and the second ones are underlined.
Backbone Method DomainNet Avg Hyper-parameters
C I P Q R S
ViT-B/16 Source (He et al., 2016) 71.62 25.59 57.34 18.07 71.90 57.24 50.29 nan
BN (Schneider et al., 2020) 0.00 0.00 0.00 0.00 0.00 0.00 0.00 nan
T3A (Iwasawa & Matsuo, 2021) 73.56 26.95 59.77 27.34 75.77 60.24 53.94 fk=100
AdaNPC (Zhang et al., 2023) 73.76 25.52 59.86 24.42 75.99 60.14 53.28 fk=200 k=5
TENT (Wang et al., 2020) 72.65 18.00 35.08 4.20 74.03 35.82 39.96 lr=1e-4
PLC (Lee, 2013) 72.29 19.27 56.69 5.00 72.45 58.04 47.29 lr=5e-5
EATA (Niu et al., 2022) 73.91 28.76 61.71 24.79 75.39 61.84 54.40 lr=1e-3
SAR (Niu et al., 2023) 73.06 17.42 40.94 11.37 73.67 40.39 42.81 lr=5e-2
TIPI (Nguyen et al., 2023) 72.71 17.62 33.37 4.54 73.98 35.96 39.70 lr=1e-4
TEA (Yuan et al., 2024) 71.96 24.17 55.31 8.83 72.20 58.00 48.41 lr=5e-5
TSD (Wang et al., 2023) 72.40 23.47 59.20 4.12 73.54 59.60 48.72 lr=1e-6 fk=50
[…]
Table 17: Accuracy comparison of different TTA methods on DomainNet dataset based on ViT-B/16 backbone. The best results are highlighted in boldface, and the second ones are underlined.
Method CIFAR-10-C Avg
@@ C I P Q R S
Table 18: Accuracy comparisons of different TTA methods on CIFAR-10-C dataset at damage level of 5, based on ResNet-18 backbone. The best results are highlighted in boldface, and the second ones are underlined.
Method CIFAR-10-C Avg
@@
Table 19: Accuracy comparisons of different TTA methods on CIFAR-10-C dataset at damage level of 5, based on ResNet-50 backbone. The best results are highlighted in boldface, and the second ones are underlined.
Method CIFAR-10-C Avg
@@
T3A (Iwasawa & Matsuo, 2021) 47.84 52.78 51.52 83.16 73.06 83.35 85.66 87.04 84.96 79.53 92.72 36.44 84.49 85.72 84.81 74.21
AdaNPC (Zhang et al., 2023) 48.70 54.14 50.95 83.31 74.10 83.46 85.86 87.20 85.31 80.09 92.70 37.92 84.70 85.84 84.70 74.60
TENT (Wang et al., 2020) 37.25 44.31 39.94 83.16 70.31 83.54 85.80 87.15 85.05 79.19 92.75 29.73 84.73 84.68 84.58 71.48
PLC (Lee, 2013) 37.18 44.27 39.84 83.15 70.31 83.54 85.77 87.15 85.07 79.19 92.75 29.72 84.74 84.69 84.59 71.46
@@
SAR (Niu et al., 2023) 37.25 44.31 39.94 83.16 70.31 83.54 85.80 87.15 85.06 79.19 92.75 29.73 84.73 84.68 84.58 71.48
TIPI (Nguyen et al., 2023) 37.24 44.32 39.93 83.17 70.30 83.55 85.77 87.16 85.07 79.18 92.74 29.73 84.75 84.69 84.58 71.48
TEA (Yuan et al., 2024) 37.23 44.31 39.92 83.17 70.30 83.56 85.79 87.15 85.06 79.20 92.75 29.73 84.74 84.69 84.58 71.48
@@
Table 20: Accuracy comparisons of different TTA methods on CIFAR-10-C dataset at damage level of 5, based on ViT-B/16 backbone. The best results are highlighted in boldface, and the second ones are underlined.
Method CIFAR-100-C Avg
@@
Table 21: Accuracy comparisons of different TTA methods on CIFAR-100-C dataset at damage level of 5, based on ResNet-18 backbone. The best results are highlighted in boldface, and the second ones are underlined.
Method CIFAR-100-C Avg
@@
Table 22: Accuracy comparisons of different TTA methods on CIFAR-100-C dataset at damage level of 5, based on ResNet-50 backbone. The best results are highlighted in boldface, and the second ones are underlined.
Method CIFAR-100-C Avg
@@
Table 23: Accuracy comparisons of different TTA methods on CIFAR-100-C dataset at damage level of 5, based on ViT-B/16 backbone. The best results are highlighted in boldface, and the second ones are underlined.
Method ImageNet-C Avg
Gau. Sho. Imp. Def. Gla. Mot. Zoo. Sno. Fro. Fog Bri. Con. Ela. Pix. Jpe.
Source (He et al., 2016) 1.54 2.27 1.48 11.44 8.68 11.12 17.62 10.64 16.21 14.02 51.52 3.44 16.49 23.35 30.67 14.70
BN (Schneider et al., 2020) 13.65 14.84 14.17 11.95 13.04 23.34 33.89 29.18 28.42 40.80 58.11 12.09 38.92 44.35 37.08 27.59
T3A (Iwasawa & Matsuo, 2021) 1.61 2.35 1.65 10.57 8.20 10.12 17.38 11.03 16.14 15.19 49.23 3.24 18.00 23.37 30.31 14.56
AdaNPC (Zhang et al., 2023) 1.42 2.01 1.42 8.23 6.49 7.64 13.82 8.50 12.08 11.97 42.81 2.77 15.41 19.92 24.49 11.93
TENT (Wang et al., 2020) 23.45 25.71 24.08 18.79 20.90 33.54 42.85 39.64 32.95 50.36 60.13 10.68 48.81 51.96 46.98 35.39
PLC (Lee, 2013) 13.64 14.85 14.16 11.96 13.02 23.36 33.91 29.18 28.43 40.78 58.12 12.08 38.91 44.35 37.08 27.59
EATA (Niu et al., 2022) 28.24 30.16 28.88 25.30 25.74 36.61 43.71 41.80 36.42 50.87 59.12 31.75 49.10 52.33 47.82 39.19
SAR (Niu et al., 2023) 28.04 29.59 27.88 23.66 23.90 36.16 43.40 40.94 36.71 51.01 60.18 27.38 48.95 52.47 47.98 38.55
TIPI (Nguyen et al., 2023) 24.45 26.52 24.75 20.37 22.25 33.65 42.46 39.31 33.47 49.93 59.44 12.53 48.41 51.51 46.92 35.73
TEA (Yuan et al., 2024) 18.82 20.50 19.00 16.27 17.68 28.51 39.17 35.19 32.26 46.92 59.16 15.42 44.39 48.81 43.64 32.38
TSD (Wang et al., 2023) 15.60 16.99 16.13 15.59 15.41 28.69 38.07 32.92 30.01 45.90 58.69 7.62 41.06 47.47 41.52 30.11
[…]
Table 24: Accuracy comparisons of different TTA methods on ImageNet-C dataset at damage level of 5, based on ResNet-18 backbone. The best results are highlighted in boldface, and the second ones are underlined.
Method ImageNet-C Avg
Gau. Sho. Imp. Def. Gla. Mot. Zoo. Sno. Fro. Fog Bri. Con. Ela. Pix. Jpe.
Source (He et al., 2016) 3.00 3.70 2.64 17.91 9.74 14.71 22.45 16.60 23.06 24.01 59.12 5.38 16.51 20.87 32.63 18.15
BN (Schneider et al., 2020) 16.32 17.09 16.97 15.23 15.54 26.64 39.38 34.46 33.45 48.43 65.67 17.08 44.62 49.49 40.47 32.06
T3A (Iwasawa & Matsuo, 2021) 2.97 3.38 2.65 17.05 9.37 13.69 22.63 16.98 22.83 25.34 57.68 5.05 18.41 20.31 32.36 18.05
AdaNPC (Zhang et al., 2023) 2.61 3.00 2.36 14.19 7.92 11.14 18.88 14.22 18.65 22.08 52.65 4.21 16.57 17.73 28.04 15.62
TENT (Wang et al., 2020) 26.70 28.90 28.01 25.01 24.72 38.66 48.84 46.04 40.49 57.04 67.96 24.42 53.73 57.72 51.54 41.32
PLC (Lee, 2013) 16.32 17.09 16.96 15.24 15.54 26.64 39.39 34.44 33.45 48.44 65.66 17.07 44.61 49.50 40.47 32.06
EATA (Niu et al., 2022) 35.38 37.75 36.34 33.29 32.74 47.51 53.07 52.53 46.30 60.54 68.06 43.82 58.42 61.15 55.72 48.17
SAR (Niu et al., 2023) 34.39 35.42 35.77 32.27 31.15 45.34 51.58 50.03 44.20 59.27 68.14 35.58 56.82 60.06 54.52 46.30
TIPI (Nguyen et al., 2023) 27.96 31.36 31.37 24.74 24.39 42.68 49.93 48.01 37.40 57.94 66.51 16.66 56.19 58.94 53.90 41.87
TEA (Yuan et al., 2024) 22.51 23.94 22.70 20.70 21.00 36.06 46.93 44.91 39.37 56.06 67.01 22.79 52.65 56.86 49.99 38.90
TSD (Wang et al., 2023) 18.90 18.80 19.63 18.31 17.49 31.12 44.27 39.74 36.02 53.38 67.00 13.06 47.97 54.54 45.95 35.08
[…]
Table 25: Accuracy comparisons of different TTA methods on ImageNet-C dataset at damage level of 5, based on ResNet-50 backbone. The best results are highlighted in boldface, and the second ones are underlined.
Method ImageNet-C Avg
Gau. Sho. Imp. Def. Gla. Mot. Zoo. Sno. Fro. Fog Bri. Con. Ela. Pix. Jpe.
Source (He et al., 2016) 35.09 32.16 35.88 31.42 25.31 39.45 31.55 24.47 30.13 54.74 64.48 48.98 34.20 53.17 56.45 39.83
T3A (Iwasawa & Matsuo, 2021) 27.87 28.15 30.27 32.42 27.00 40.66 33.52 25.95 30.76 56.32 64.85 50.20 37.99 53.75 57.01 39.78
AdaNPC (Zhang et al., 2023) 30.01 26.86 30.98 28.19 23.40 36.38 29.65 21.18 26.59 52.78 61.24 44.53 34.44 50.50 54.98 36.78
TENT (Wang et al., 2020) 51.19 50.00 52.48 47.35 42.95 54.40 45.19 7.43 16.27 64.87 70.90 64.35 25.97 63.35 63.48 48.01
PLC (Lee, 2013) 29.10 29.61 31.58 31.19 25.10 39.33 31.45 24.58 30.18 54.33 64.47 48.44 34.17 52.52 55.11 38.74
EATA (Niu et al., 2022) 56.63 56.20 57.49 56.13 57.13 62.26 62.89 64.02 62.77 73.71 77.07 70.34 67.39 71.42 69.96 64.36
SAR (Niu et al., 2023) 54.90 55.82 56.68 55.94 55.61 62.47 58.11 17.20 34.16 71.85 77.14 63.48 65.75 71.45 68.52 57.94
TIPI (Nguyen et al., 2023) 56.76 56.70 58.45 55.90 56.10 61.75 19.67 1.88 4.51 64.18 75.91 69.24 6.37 70.05 69.97 48.50
TEA (Yuan et al., 2024) 39.46 38.72 41.90 24.47 28.03 42.01 33.46 13.43 33.46 53.89 66.09 58.66 34.95 55.65 56.41 41.37
TSD (Wang et al., 2023) 36.78 33.68 37.61 32.28 26.36 40.93 32.65 25.20 31.49 56.17 65.65 54.27 35.10 54.61 57.17 41.33
[…]
Table 26: Accuracy comparisons of different TTA methods on ImageNet-C dataset at damage level of 5, based on ViT-B/16 backbone. The best results are highlighted in boldface, and the second ones are underlined.